OpenStudy (anonymous):

Which value of x makes the following equation true? 3(x - 3) - 14 = 2(x - 7) + 10 A. 16 B. 26 C. 12 D. 19

jigglypuff314 (jigglypuff314):

the answer is D. 19 because when you distribute 3(x - 3) - 14 = 2(x - 7) + 10 it becomes 3x - 9 - 14 = 2x - 14 +10 which can be reduced to 3x - 23 = 2x -4 +23 +23 3x = 2x +19 -2x -2x x = 19